# Complete Chess Games Enable LLM Become A Chess Master

### v1

@@ Abstract
Yinqi Zhang, Xintian Han, Haolong Li, Kedi Chen, Shaohui Linthanks: Corresponding Author,
@@ 1 Introduction
Recently, Large Language Models (LLMs) based on transformer architectures Vaswani et al. (2017) have demonstrated capabilities well beyond language modeling. A key milestone was the advent of ChatGPT Ouyang et al. (2022). Extensive research has focused on developing efficient LLM base models Du et al. (2021); Biderman et al. (2023); Black et al. (2022); Computer (2023); Touvron et al. (2023), including supervised models Taori et al. (2023); Chiang et al. (2023); Anand et al. (2023); Köpf et al. (2023) and models using Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2017); Ouyang et al. (2022); Rando and Tramèr (2023); Bai et al. (2023). Recent research Wei et al. (2022); Li et al. (2024) shows that as models scale, their capabilities increase. This raises questions about LLMs’ intelligence and learning structures. Chess, an ancient game, has dialogue-like characteristics in its notational structures such as Forsyth-Edwards Notation (FEN), Standard Algebraic Notation (SAN), and Universal Chess Interface (UCI). Machine learning in chess has evolved to include reinforcement learning and neural networks based on supervised learning from human gameplay. Developments include AI-based engines like Leela Chess Zero (LC0) [^1] https://lczero.org and Stockfish NNUE [^2] https://stockfishchess.org, which refine their algorithms through new learning. Deep learning has shown the potential of AI in strategic games. The ChessGPT Feng et al. (2023) model demonstrated the ability to choose optimal moves by learning from human language and chess data. However, models like ChessGPT cannot generate the best move based on the current game state and complete an entire match. Our focus is on match completeness and quality of gameplay.
@@ 2 A Large Scale Dataset of Chess
Our Best Move dataset was created through a search method using Stockfish. It consists of two parts: the short round dataset from Chessdb [^3] http://chessdb.sourceforge.net and the long round dataset from self-play endgames based on Stockfish evaluations. Stockfish, using heuristic functions and an alpha-beta game tree search, evaluates positions. We searched for valid moves from current positions, with search depths of 12-50 for short rounds and 50-200 for long rounds, limiting each search to two seconds. The highest win-rate moves were selected as the best moves.
@@ 3 Model
The Generative Pre-trained Transformer (GPT-3) is an autoregressive language model that generates human-like text through deep learning. It trains on casual language modeling, predicting the next word based on previous words. We trained a GPT-like model using open-llama-3B Geng and Liu (2023) and the chess resources from Section 2. Unlike policy behavior data in robotics or gaming, chess state and move data can be expressed textually. This allows chess to be rendered as a text-based game, enabling imitation learning for policy through casual language modeling of the game dataset (Figure LABEL:fig:datasample). This innovative approach of applying language modeling to chess signifies a novel shift in policy learning, leveraging the game’s unique aspects to develop superior gameplay tactics.
@@ 4.1 Actual Games
where $Elo_{N}$ is the updated Elo rating after the game. $Elo_{O}$ is the previous Elo rating before the game. $K$ is the weight of the tournament. In professional chess, $K$ is often set to 10 for high-ranked players and 20 for low-ranked players. $R_{A}$ is the actual result of the game (1 for win, 0.5 for draw, 0 for loss). $R_{E}$ is the expected result of the game. $Elo_{S}$ is the old Elo rating of Stockfish. $Elo_{M}$ is the old Elo rating of the model.
Moreover, we refer to the method introduced by Stockfish to convert between its skill level and Elo rating. The specific calculation method is shown as follows.
@@ 5 Experiment Analysis
5.1 Evaluation Set
Stockfish ChessLLM
Skill level Elo Win Lose Draw Elo
ChessLLM 0 1350-1440 61 29 10 1632 $\pm$ 45
vs. 1 1450-1560 56 37 7 1753 $\pm$ 55
Stockfish 2 1570-1720 30 69 1 1788 $\pm$ 75
@@ 5.4 Compare with Other LMs
We conduct matches between ChessLLM and other Language Models (LMs) such as LLAMA Touvron et al. (2023), RedPajama Computer (2023), ChessGPT-Base Feng et al. (2023), and ChessGPT-Chat Feng et al. (2023), calculating their respective win rating. As other models cannot guarantee the legality of the moves they generate, we bring in Stockfish to aid in this process. Should the model fail to produce a valid move even after 50 sampling efforts, a mechanism is employed wherein there’s a 50% chance of favoring either the best move identified by Stockfish or a randomly picked move from the list of all possible legal moves. Similarly, as ChessGPT is unable to generate the best move for the next step, we generate all legal moves through Stockfish and utilize their proposed general policy for selection, picking the most optimal move as recognized by the model.

### v2

@@ Abstract
Yinqi Zhang, Xintian Han, Haolong Li, Kedi Chen, Shaohui Linthanks: Corresponding Author
@@ 1 Introduction
Recently, Large Language Models (LLMs) based on transformer architectures Vaswani et al. (2017) have demonstrated capabilities well beyond language modeling. A key milestone was the advent of ChatGPT Ouyang et al. (2022). Extensive research has focused on developing efficient LLM base models Du et al. (2021); Biderman et al. (2023); Black et al. (2022); Computer (2023); Touvron et al. (2023a), including supervised models Taori et al. (2023a); Chiang et al. (2023); Anand et al. (2023); Köpf et al. (2023) and models using Reinforcement Learning from Human Feedback (RLHF) Christiano et al. (2017); Ouyang et al. (2022); Rando and Tramèr (2023); Bai et al. (2023). Recent research Wei et al. (2022); Li et al. (2024) shows that as models scale, their capabilities increase. This raises questions about LLMs’ intelligence and learning structures. Chess, an ancient game, has dialogue-like characteristics in its notational structures such as Forsyth-Edwards Notation (FEN), Standard Algebraic Notation (SAN), and Universal Chess Interface (UCI). Machine learning in chess has evolved to include reinforcement learning and neural networks based on supervised learning from human gameplay. Developments include AI-based engines like Leela Chess Zero (LC0) [^1] https://lczero.org and Stockfish NNUE [^2] https://stockfishchess.org, which refine their algorithms through new learning. Deep learning has shown the potential of AI in strategic games. The ChessGPT Feng et al. (2023) model demonstrated the ability to choose optimal moves by learning from human language and chess data. However, models like ChessGPT cannot generate the best move based on the current game state and complete an entire match. Our focus is on match completeness and quality of gameplay.
@@
2 Related work
2.1 Large Language Model
The emergence of Large language models (LLMs) GPT-4 Achiam et al. (2023), stands as a noteworthy testament to the significant advancements in natural language understanding and generation. Unlike commercial models, open-source models, such as Alpaca Taori et al. (2023b), Vicuna Zheng et al. (2023), and Llama2 Touvron et al. (2023b), have recently become more accessible. Due to their proficiency in text reasoning, LLMs are increasingly being utilized in everyday applications(Chen et al., 2024). Comprehensive benchmarks, such as MMLU Hendrycks et al. (2021) and HELM Lee et al. (2023), have been devised for thorough assessments of the LLMs’ overall capabilities. Our work takes this evaluation process one step further, particularly highlighting and investigating the capacity of LLMs’ ability to play abstract games.
2.2 Supervised Fine-tuning
Supervised Fine-tuning has emerged as a revolutionary technique within the field of machine learning and has been the subject of a multitude of studies. Owing to the continuous advancements in the domain of transfer learning, pre-trained models, fine-tuned in a supervised manner, have demonstrated superior performance in numerous tasks. Notably, in the context of natural language processing (NLP), the work by Howard and Ruder became a pioneering model of this technique. Their method Howard and Ruder (2018) leverages the power of transfer learning for comprehensive language modeling tasks, thus effectively surpassing previous benchmarks. Manipulating the same concept, BERT Devlin et al. (2019), an innovative model fine-tuned in a supervised manner for a wide array of NLP tasks. BERT demonstrated remarkable success within various NLP tasks, setting new performance standards.
In this work, we trained ChessLLM with supervised fine-tuning.
2.3 Chess
The quest to develop artificial intelligence capable of playing chess can be traced back to the inception of computer science Turing (1953); Campbell et al. (2002). The application of machine learning, particularly deep learning, in the domain of chess has been explored extensively in recent years Silver et al. (2018); McGrath et al. (2022). One of the pivotal works in this field is the study by DeepChess David et al. (2016), which presented an end-to-end learning method for chess based solely on deep neural networks, demonstrating the powerful capabilities of machine learning in comprehending and mastering strategic games without a priori knowledge.
In this work, we applied LLMs to chess and evaluated them with Elo rating.
@@ 3 A Large Scale Dataset of Chess
Our Best Move dataset was created through a search method using Stockfish. It consists of two parts: the short round dataset from Chessdb [^3] http://chessdb.sourceforge.net and the long round dataset from self-play endgames based on Stockfish evaluations. Stockfish evaluates positions using heuristic functions and an alpha-beta game tree search. We searched for valid moves from current positions, with search depths of 12-50 for short rounds and 50-200 for long rounds, limiting each search to two seconds. The highest win-rate moves were selected as the best moves.
@@ 4 Model
The Generative Pre-trained Transformer (GPT-3) is an autoregressive language model that generates human-like text through deep learning. It trains on casual language modeling, predicting the next word based on previous words. We trained a GPT-like model using open-llama-3B Geng and Liu (2023) and the chess resources from Section 3. Unlike policy behavior data in robotics or gaming, chess state and move data can be expressed textually. This allows chess to be rendered as a text-based game, enabling imitation learning for policy through casual language modeling of the game dataset (Figure 2). This innovative approach of applying language modeling to chess signifies a novel shift in policy learning, leveraging the game’s unique aspects to develop superior gameplay tactics.
Figure 2: One example of training data.
@@ 5.1 Actual Games
where $Elo_{N}$ is the updated Elo rating after the game. $Elo_{O}$ is the previous Elo rating before the game. $K$ is the weight of the tournament. In professional chess, $K$ is often set to 10 for high-ranked players and 20 for low-ranked players. $R_{A}$ is the actual result of the game (1 for win, 0.5 for draw, 0 for loss). $R_{E}$ is the expected result of the game. $Elo_{S}$ is the old Elo rating of Stockfish. $Elo_{M}$ is the old Elo rating of the model. Moreover, we refer to the method introduced by Stockfish to convert between its skill level and Elo rating. The specific calculation method is shown as follows.
@@ 6 Experiment Analysis
6.1 Evaluation Set
Stockfish ChessLLM
Skill level Elo Win Lose Draw Elo
ChessLLM 0 1350-1440 61 29 10 1632 $\pm$ 45
vs. 1 1450-1560 56 37 7 1753 $\pm$ 55
Stockfish 2 1570-1720 30 69 1 1788 $\pm$ 75
@@ 6.4 Compare with Other LMs
We conduct matches between ChessLLM and other Language Models (LMs) such as LLAMA Touvron et al. (2023a), RedPajama Computer (2023), ChessGPT-Base Feng et al. (2023), and ChessGPT-Chat Feng et al. (2023), calculating their respective win rating. As other models cannot guarantee the legality of the moves they generate, we bring in Stockfish to aid in this process. Should the model fail to produce a valid move even after 50 sampling efforts, a mechanism is employed wherein there’s a 50% chance of favoring either the best move identified by Stockfish or a randomly picked move from the list of all possible legal moves. Similarly, as ChessGPT is unable to generate the best move for the next step, we generate all legal moves through Stockfish and utilize their proposed general policy for selection, picking the most optimal move as recognized by the model.